OpenStudy (precal):

Find the intervals on which f is increasing and decreasing. f(x)=cos^2(x) on [-pi, pi]

OpenStudy (bahrom7893):

Increasing when y'>0, decreasing when y'<0

OpenStudy (precal):

\[f(x)=\cos^2(x)\] \[f ' (x)=-2\sin(x)\cos(x)\]

OpenStudy (precal):

Is that using the derivative? I have to show sign analysis

OpenStudy (bahrom7893):

Set the derivative equal to 0 and solve.

OpenStudy (bahrom7893):

Then you'll get some values suppose when y'=0, x1=a, x2=b, x3=c. Create something like a table like the one I'll show you, very helpful...

OpenStudy (precal):

Am I suppose to do -2sin(x)cos(x)>0 to find where it is increasing?

OpenStudy (bahrom7893):

 By the way, those are just made up functions.